OpenStudy (anonymous):

Which formula could not be used to calculate the volume of this square-based pyramid? v=1/3Bh v=1/3/wh v=1/3s2h v=pi2h Which formula could not be used to calculate the volume of this square-based pyramid? v=1/3Bh v=1/3/wh v=1/3s2h v=pi2h @Mathematics

OpenStudy (anonymous):

please help me

OpenStudy (across):

The first three choices have a 1/3 in them. The answer looks like it's the last one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but the last one is v=1/3 pi2h

OpenStudy (anonymous):

so they all have 1/3

OpenStudy (anonymous):

The volume of a square-base pyramid is: \[\Large V = \frac{1}{3}{b^2}h\] Where b is any one of the side lengths of the pyramid's base (since it's square, they're all the same), and h is the height. So assuming: \[\Large \begin{array}{l} {\rm{v}} = {\rm{1}}/{\rm{3s2h}}\\ = {\rm{v}} = {\rm{1}}/{\rm{3}}{{\rm{s}}^{\rm{2}}}{\rm{h}} \end{array}\] Where s is a side length, then that's the formula.

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

Which one is the answer?